A piece of lead falls from a height of 100 meters on a fixed non _ conducting slad which brings it to rest. Show that it is temperature immediately 7.5°c, knowing that the specific heat of lead is 0.13 J/g°c.)


1
Expert's answer
2021-05-17T18:50:39-0400

When the lead ball falls on the non-conducting surface, the lost potential energy is absorbed by the ball as heat energy.

Potential energy of lead "=" thermal energy of slab


"mgh=Cm\varDelta T"

"\varDelta T=\dfrac{gh}{C}"

Given "g=9.8\ m/s^{2}, h=100\ m,"

"C=0.13\ J/(g \degree C)=130\ J/(kg \degree C)"

Substitute


"\varDelta T=\dfrac{9.8\ m/s^{2}\cdot100\ m}{130\ J/(kg \degree C)}=\dfrac{98}{13}\ \degree C"

"\varDelta T=7.5\ \degree C"